how do u do this 2x+3y=(-4)
x+y=(-2)

and this x+2y=0
y=3

Answer by Paul(988)

x+y=-2 (equation 1) Mulitply equation 1 by 3)
2x+3y=-4 (equation 2)

=3x-3y=-6
2x-3y=-4
x=-2